# Morphology-Specific Peptide Discovery via Masked Conditional Generative Modeling

### v2

@@ Abstract
Peptide self-assembly prediction offers a powerful bottom-up strategy for designing biocompatible, low-toxicity materials for large-scale synthesis in a broad range of biomedical and energy applications. However, screening the vast sequence space for categorization of aggregate morphology remains intractable. We introduce PepMorph, an end-to-end peptide discovery pipeline that generates novel sequences that are not only prone to aggregate but self-assemble into a specified fibrillar or spherical morphology. We compiled a new dataset by leveraging existing aggregation propensity datasets and extracting geometric and physicochemical isolated peptide descriptors that act as proxies for aggregate morphology. This dataset is then used to train a Transformer-based Conditional Variational Autoencoder with a masking mechanism, which generates novel peptides under arbitrary conditioning. After filtering to ensure design specifications and validation of generated sequences through coarse-grained molecular dynamics simulations, PepMorph yielded 83% accuracy in intended morphology generation, showcasing its promise as a framework for application-driven peptide discovery.
Introduction
@@
Supramolecular self-assembly is a powerful bottom-up strategy for designing functional materials. Small molecular building blocks can spontaneously organize into well-ordered architectures through weak non-covalent interactions (hydrogen bonding, aromatic $\pi$-$\pi$ stacking, hydrophobic effects, electrostatics, metal coordination, etc.). Given the dynamic and reversible nature of these interactions, supramolecular assemblies often exhibit adaptive, self-healing, and stimuli-responsive behaviors [1]. Such processes are ubiquitous in nature; for example, phospholipid molecules assemble into cell membranes, and complementary strands of nucleotides form the DNA double helix via hydrogen bonding. These biological motifs have inspired a variety of synthetic supramolecular materials.
@@
Among the various supramolecular building blocks, peptides stand out for their inherent biocompatibility, chemical tunability, and straightforward synthesis  [2, 3]. The combinatorial diversity afforded by the 20 canonical amino acids defines a vast possible design space for peptide assemblies, as even minor modifications in sequence often yield markedly different supramolecular outcomes, highlighting the critical role of sequence design in defining material structure and function. Indeed, different peptide sequences can form a rich variety of nanostructures — including fibers, tubes, rods, sheets, vesicles, and micelles —, depending on their primary sequence [4]. Moreover, environmental factors such as pH, temperature, and solvent composition can also steer the assembly trajectory and thereby dictate the final morphology  [5].
@@
Peptide-based supramolecular materials have found utility across a diverse array of applications. In biomedicine, for example, peptide assemblies have been employed as drug-delivery vehicles, tissue-engineering scaffolds, biosensors, and theranostic agents  [6, 7, 8, 9, 10]. These examples emphasize the potential of peptide self-assembly in life-science and materials-science contexts. However, realizing this potential depends on discovering sequences that adopt target morphologies and satisfy functional requirements, a highly complex challenge given the vast combinatorial sequence space.
@@
Computational design strategies have therefore become necessary for exploring this sequence space. While trial-and-error synthesis is impractical for large-scale screening, conventional rational design, relying on expert intuition and existing heuristics, can be biased and may fail to capture unexpected solutions. In the past few years, machine learning has emerged as a powerful alternative for peptide sequence design, particularly through generative models. Variational autoencoders (VAEs) [11] learn continuous, low-dimensional embeddings of sequences, while conditional VAEs (CVAEs) [12] further enable targeted generation by incorporating property labels or descriptors as conditioning inputs. For instance, PepCVAE demonstrated semi-supervised generation of antimicrobial peptides by conditioning on activity labels [13], as well as the HydrAMP framework built upon this approach to optimize peptide potency and hydrophobic balance for antimicrobial activity [14].
Although developed for images or general tabular data rather than for peptide design, recent work has also shown that CVAEs can condition on an arbitrary subset of attributes by explicitly modeling a mask of observed covariates. In VAEAC (Variational Autoencoder with Arbitrary Conditioning), a conditional prior $p(z\mid c,m)$ processes the observed conditions $c$ and a binary mask $m$. In contrast, the decoder is conditioned on ($c$,$m$) and trained under random masks to handle any subset at test time [15]. Related approaches extend CVAEs to missing-covariate settings by marginalizing unobserved conditions during training [16]. These approaches highlight the ability of CVAE-based frameworks to efficiently explore complex design spaces and generate samples conditioned on fully specified or arbitrary partial sets of attributes, laying the groundwork for extending generative design to peptide self-assembly and morphology control, given that the labeled search space is still not extensive.
Indeed, peptide study via machine learning models is only beginning to take hold when applied to self-assembly design, especially given the lack of data. Recently, larger-scale in silico screening of the peptide sequence space has enabled the creation of the first extensive aggregation-propensity datasets [17, 18, 19]. Aggregation Propensity (AP) quantifies the ratio of solvent-accessible surface area (SASA) at the end versus the beginning of a Molecular Dynamics (MD) trajectory [20]. This metric is a proxy for a peptide’s inherent tendency to self-associate and has been used to build classifiers for aggregation given solely the peptide sequences in FASTA format [18]. Despite recent progress, peptide generation for self-assembly remains largely unexplored. Njirjak et al. tackle this problem by first training a supervised recurrent neural network classifier on a curated set of 368 peptides and then using it as a fitness oracle within a genetic algorithm to propose sequences with high predicted self-assembly propensity [21]. Candidates are validated by coarse-grained MD (CG-MD) and, for a subset, experimentally, yielding a reported discovery accuracy of $80-95\%$. However, genetic algorithm search explores only a narrow region of sequence space and produces high redundancy — generated sequences show $40\%$ similarity — limiting diversity and coverage of the peptide aggregate space.
@@
Beyond these limitations, morphology remains a critical determinant of material function even when aggregation-prone sequences are identified. Applications ranging from drug delivery to biosensors demand specific aggregate geometries (fibers, tubes, spheres, etc.), yet datasets linking peptide sequence to computationally or experimentally validated self-assembly morphology are scarce. To date, efforts for data gathering concerning morphology have primarily assembled experimental databases such as SAPdb [22]; however, these are small and very sparse datasets, as simulation setups differ substantially across the aggregated studies. This lack of scale and consistency renders the available data unsuitable for training machine learning models. Without direct morphology labels, generative design must rely on indirect proxies as stand-ins for the desired assembly shape; for example, predicted descriptors derived from the peptide sequence and the isolated peptide structure model. Features such as $\beta$-sheet formation, amphiphilicity, and charge distribution were postulated to directly influence whether peptides self-assemble into fibrils, sheets, or globular micelles [23]. For example, sequences dominated by hydrophobic residues may form compact, spherical aggregates (minimizing exposed hydrophobic surface). Tools like PEP-FOLD [24] allow the effective generation of the most likely 3D conformations of isolated peptides in aqueous solution, which is one of the most common target solvents for supramolecular peptide assemblies.
@@
Computationally derived peptide-level descriptors may, therefore, effectively predict a peptide’s assembly behavior. To validate this hypothesis, we introduce PepMorph, a framework designed to guide peptide discovery through aggregation propensity and morphology awareness. At its core, PepMorph employs a transformer-based CVAE with a masking mechanism, akin to VAEAC, allowing peptide generation to be flexibly conditioned on descriptors of isolated peptides that serve as morphology proxies. We demonstrate that the model generates highly novel and diverse peptides, with $10\%$ similarity, whose properties match all their targets in $55\%$ of cases. All the generated candidates are then passed through a dual-stage filtering pipeline, ensuring that only sequences meeting aggregation and sequence or structure requirements are retained. In end-to-end validation with CG-MD simulations, all sequences formed aggregates, and $83\%$ matched the intended morphology by visual inspection, supporting morphology-aware sequence discovery. PepMorph bridges the gap between unguided sequence exploration and fully constrained design, offering a versatile route toward developing functional supramolecular materials.
Results
@@
We begin by forming the PepMorph dataset, an extensive aggregation and isolated peptide descriptor dataset. Then, we implement and rigorously evaluate the PepMorph pipeline on tri-to-decapeptide morphology-aware aggregate discovery tasks. To do so, we train a transformer-based CVAE with a masking mechanism on an augmented dataset comprising aggregation metrics and computed descriptors for each peptide sequence. We then demonstrate the model’s generative performance by sampling sequences conditioned on assembly and structural metrics. Next, we apply our post-generation filters, selecting fifteen candidates each for fibrillar and spherical morphologies. Finally, we subject these peptides to CG-MD simulations in aqueous solution, quantifying their AP and examining the resulting aggregate morphologies quantitatively and visually.
@@
One of the main issues with predictive models for short peptides is the lack of extensive datasets like the Protein Data Bank, which provides a comprehensive and standardized structural reference for proteins [25]. Nevertheless, we can move towards machine learning models for assembly of short peptides by enriching existing datasets with additional sequences and morphology proxy features (Figure 1a).
@@
Figure 1: PepMorph dataset (a) Data curation and feature-extraction workflow: we merge three sources: Wang et al. [17] ($\sim$62k peptides, 5-10 amino acids ($a$)), Teijlingen & Tuttle [19] ($\sim$60k, 3-8 $a$), and a set of $\sim$39k random peptides (4-10 $a$). After deduplication, self-assembly (SA) labels are assigned, and peptide conformations are predicted with PEP-FOLD to derive all descriptors (sequence length, $\beta$-sheet content, net charge and hydrophobic moment). The resulting PepMorph corpus contains  161k unique peptides spanning 3-10 $a$ with aggregation-propensity (AP) values and SA/no-SA labels, as well the calculated peptide–level descriptors. Univariate summaries of the PepMorph dataset are shown, specifically of AP density (b), assembly vs no assembly (c), hydrophobic moment (d), peptide length (e), presence of $\beta$-sheet content (f) and net-charge (g). Regions regarding no-assembly and assembly are highlighted in (b), and condition regions used when targeting specific morphologies are highlighted in the remaining summaries (d-g).
@@
We curated continuous AP values from Wang et al. [17] and Teijlingen & Tuttle [19], where AP follows the CG-MD-derived SASA ratio [17]. Both studies used the MARTINI 2.2 force field. For the 56 peptides common to both datasets, we observed a mean AP difference of $0.089\pm 0.13$, supporting comparability, and thus mergeability, of the two AP sources. During deduplication, we retained the values reported by Wang et al. and, following their procedure, subsequently assigned peptide a self-assembly/no-self-assembly (SA/no-SA) flag based on AP: $\geq\!1.8$ indicates aggregation, $\leq\!1.65$ indicates no aggregation, and values in between remain undefined (Figure 1b) [17]. This yielded a merged dataset with AP for $121,652$ peptides, $93,668$ with a SA/no-SA label.
Given the masking mechanism of our model, further explained in the next section, PepMorph does not require any specific target to be present to generate sequences (e.g., AP can be absent). As such, we also added $39,468$ sequences not present in the literature–derived set to expand the conditioning space. For these and for the set of peptides from Wang et al.’s work, we further processed the peptides through PEP-FOLD [24] to generate isolated 3D conformations and extract, for each, key properties that may drive specific assemblies. As we will elaborate in the validation pipeline, we focus on three biophysical metrics that plausibly steer spherical versus fibrillar assembly:
@@
  • $\beta$-sheet content, referring to whether or not any of the amino acid residues adopt a $\beta$-strand conformation; peptides with $\beta$-sheet propensity tend to stack via backbone hydrogen bonds into elongated, fibrillar (cylindrical) assemblies [26].
@@
  • net charge, measuring the peptide’s overall charge at physiological pH; for example, highly charged peptides experience strong electrostatic repulsion that impedes self-assembly [26].
We obtained valid 3D conformations (and thus retrievable descriptors) for $80,870$ sequences ($\approx 50\%$). The resulting corpus contains $161,120$ unique peptide records with their composition described in Table 1. The distribution across sequence lengths is relatively balanced, with between $15,000$ and $35,000$ peptides represented at each length (Figure 1e). All possible tripeptides are included, but for longer lengths the dataset covers a smaller fraction of the exponentially expanding theoretical space of $20^{L}$ sequences (see Supplementary Material).
Table 1: PepMorph dataset composition ($161,120$ entries), comprising the peptide sequence and length, two aggregation descriptors, and three morphology–proxy descriptors. Coverage counts non-null entries.
@@
sequence Peptide in FASTA format (uppercase single-letter codes) [27]. [17, 19] + random $161,120$ ($100.0\%$)
length Number of residues in sequence. derived
ap Aggregation propensity (AP) [17, 19] $121,652$ ($75.5\%$)
is_assembled Self-assembly label (1=SA, 0=no-SA). $93,668$ ($58.1\%$)
has_beta_sheet_content Whether any residue is assigned to $\beta$-strand in the predicted peptide 3D conformation. computed $80,870$ ($50.2\%$)
hydrophobic_moment Magnitude of hydrophobic moment from predicted peptide 3D conformation.
net_charge Peptide net charge at neutral pH.
The resulting PepMorph dataset also exhibits notable imbalances across descriptors. In particular, SA labels are roughly double that of no-SA (Figure 1c), and $\beta$-sheet content is exceedingly rare (Figure 1f). The AP distribution itself is continuous, with many sequences clustering around the cutoff of $\geq\!1.8$ (Figure 1b); as such, this threshold should be regarded as an approximate separator. By contrast, continuous descriptors such as the hydrophobic moment (Figure 1d) and net charge (Figure 1g) show a broad spread. However, most peptides lie in the low-to-moderate regime for the former. These patterns emphasize the potential bias and structural diversity in the PepMorph dataset. Its associated sparsity is, however, naturally handled by the aforementioned masking mechanism. Furthermore, extending the dataset with additional descriptors from the generated peptide conformations would also be trivial to integrate into the pipeline.
@@
To address this limitation and fully leverage our dataset, we develop a transformer-based CVAE that supports partial conditioning via an explicit masking mechanism, similarly to the arbitrary–conditioning paradigm of VAEAC — although in VAEAC this is used for generic feature imputation and image inpainting, rather than targeted masking of descriptors [15] (Figure 2a).
Figure 2: PepMorph model and generation validation. (a) Schematic of the Transformer-based Conditional Variational Autoencoder with the masking mechanism: a descriptor vector $c$ and mask $m$ are summarized into a condition summary that conditions both the latent prior and the autoregressive Transformer decoder, enabling generation under arbitrary subsets of constraints. (b) Amino acid frequency in generated peptides closely follows the training distribution for both common and rare condition sets. (c) Novelty relative to the training set, quantified by the nearest-neighbour normalized edit distance (NED), showing similar novelty ability in common and rare conditions. (d) Condition-matching as a function of the number of conditioned descriptors $k$: the fraction of peptides meeting their targets declines as constraints tighten. (e) Similarity via Needleman-Wunsch percent identity of generated sequences (points) to the training set ($\mathrm{Sim}_{\text{train}}$) vs. generated sequences within the same common conditions set ($\mathrm{Sim}_{\text{gen}}^{\text{within}}$), color coded by the number of conditions $k$; values remain near low-identity baselines. (f) Summary table of generation metrics: high exact novelty, broad diversity, low global similarity, and strong condition fidelity.

### v3

# Morphology-Aware Peptide Discovery via Masked Conditional Generative Modeling

@@ Abstract
Peptide self-assembly prediction offers a powerful bottom-up strategy for designing biocompatible, low-toxicity materials for large-scale synthesis in a broad range of biomedical and energy applications. However, screening the vast sequence space for categorization of aggregate morphology remains intractable. We introduce PepMorph, an end-to-end peptide discovery pipeline that generates novel sequences that are not only prone to aggregate but whose self-assembly is steered toward fibrillar or spherical morphologies by conditioning on isolated peptide descriptors that serve as morphology proxies. To this end, we compiled a new dataset by leveraging existing aggregation propensity datasets and extracting geometric and physicochemical descriptors. This dataset is then used to train a Transformer-based Conditional Variational Autoencoder with a masking mechanism, which generates novel peptides under arbitrary conditioning. After filtering to ensure design specifications and validation of generated sequences through coarse-grained molecular dynamics (CG-MD) simulations, PepMorph yielded 83% success rate under our CG-MD validation protocol and morphology criterion for the targeted class, showcasing its promise as a framework for application-driven peptide discovery.
keywords:
peptide self-assembly, peptide discovery, aggregate morphology control, conditional variational autoencoder, arbitrary conditioning, aggregation propensity
[…]
{tocentry}
1 Introduction
@@
Supramolecular self-assembly is a powerful bottom-up strategy for designing functional materials. Small molecular building blocks can spontaneously organize into well-ordered architectures through weak non-covalent interactions (hydrogen bonding, aromatic $\pi$-$\pi$ stacking, hydrophobic effects, electrostatics, metal coordination, etc.). Given the dynamic and reversible nature of these interactions, supramolecular assemblies often exhibit adaptive, self-healing, and stimuli-responsive behaviors aidaFunctionalSupramolecularPolymers2012. These principles have inspired a variety of synthetic supramolecular materials whose structure and function emerge from collective, non-covalent organization.
@@
Among the various supramolecular building blocks, peptides stand out for their inherent biocompatibility, chemical tunability, and straightforward synthesis okesolaMulticomponentSelfassemblyTool2018a, sheehanPeptideBasedSupramolecularSystems2021. The combinatorial diversity afforded by the 20 canonical amino acids defines a vast possible design space for peptide assemblies, as even minor modifications in sequence often yield markedly different supramolecular outcomes, highlighting the critical role of sequence design in defining material structure and function. Indeed, different peptide sequences can form a rich variety of nanostructures — including fibers, tubes, rods, sheets, vesicles, and micelles — depending on their primary sequence makamMinimalisticPeptideSupramolecular2018a. At the same time, supramolecular morphology is not purely sequence-intrinsic: assembly outcomes are often co-determined by experimental conditions (e.g., concentration, solvent, pH/ionic strength, temperature) and preparation history, such that the same sequence may yield different morphologies under different contexts wehnerSupramolecularPolymerizationKinetic2020, kaygisizContextDependenceAssembly2025a, drazicUsingMachineLearning2025b.
@@
Peptide-based supramolecular materials have found utility across a diverse array of applications. In biomedicine, for example, peptide assemblies have been employed as drug-delivery vehicles, tissue-engineering scaffolds, biosensors, and theranostic agents  guptaUltrashortPeptideSelfAssembly2020a, liPeptidemodulatedSelfassemblyVersatile2019, ashworthPlasticsProteins2022, boddulaPeptidebasedNovelSmall2021, nguyenSelfassemblyDiphenylalaninePeptide2016. These examples emphasize the potential of peptide self-assembly in life-science and materials-science contexts. However, realizing this potential depends on discovering sequences that adopt target morphologies and satisfy functional requirements, a highly complex challenge given the vast combinatorial sequence space.
@@
Computational design strategies have therefore become necessary for exploring this sequence space. While trial-and-error synthesis is impractical for large-scale screening, conventional rational design, relying on expert intuition and existing heuristics, can be biased and may fail to capture unexpected solutions. In the past few years, machine learning has emerged as a powerful alternative for peptide sequence design, particularly through generative models. Variational autoencoders (VAEs) kingmaAutoEncodingVariationalBayes2022 learn continuous, low-dimensional embeddings of sequences, while conditional VAEs (CVAEs) sohnLearningStructuredOutput2015 further enable targeted generation by incorporating property labels or descriptors as conditioning inputs. For instance, PepCVAE demonstrated semi-supervised generation of antimicrobial peptides by conditioning on activity labels dasPepCVAESemiSupervisedTargeted2018, as well as the HydrAMP framework built upon this approach to optimize peptide potency and hydrophobic balance for antimicrobial activity szymczakDiscoveringHighlyPotent2023. Beyond fixed conditioning, recent CVAE formulations explicitly support arbitrary-subset conditioning by augmenting the conditioning inputs with a binary mask of observed attributes and training the decoder to remain valid when only a subset is provided (e.g., VAEAC) ivanovVariationalAutoencoderArbitrary2018. Related approaches address missing-covariate settings by marginalizing unobserved conditions during training ramchandranLearningConditionalVariational2024.
As it pertains to peptide self-assembly, machine learning has also advanced rapidly, spanning large-scale predictors, autonomous search, and human/experiment-in-the-loop discovery; however, progress remains constrained by data heterogeneity and limited standardized condition-annotated labels drazicUsingMachineLearning2025b, batraMachineLearningOvercomes2022, talluriDiscoveryUnconventionalNonintuitive2025, xuAcceleratingPredictionDiscovery2023, yangLearningRulesPeptide2025, wangDeepLearningEmpowers2023a. Recently, larger-scale in silico screening of the peptide sequence space has enabled the creation of the first extensive aggregation-propensity datasets wangDeepLearningEmpowers2023a, liuEfficientPredictionPeptide2023a, vanteijlingenTripeptidesTwoStepActive2021a. Aggregation Propensity (AP) quantifies the ratio of solvent-accessible surface area (SASA) at the end versus the beginning of a Molecular Dynamics (MD) trajectory frederixExploringSequenceSpace2015a. This metric is a proxy for a peptide’s inherent tendency to self-associate and has been used to build classifiers for aggregation given solely the peptide sequences in FASTA format liuEfficientPredictionPeptide2023a.
Building on these predictors, computational discovery pipelines can propose new sequences and validate them using simulations and/or experiments. Njirjak et al. tackle this problem by first training a supervised recurrent neural network classifier on a curated set of 368 peptides and then using it as a fitness oracle within a genetic algorithm to propose sequences with high predicted self-assembly propensity njirjakReshapingDiscoverySelfassembling2024a. Candidates are validated by coarse-grained MD (CG-MD) and, for a subset, experimentally, yielding a reported discovery accuracy of $80-95\%$. However, as with many metaheuristic search approaches, exploration can trade off diversity against exploitation of high-scoring motifs. For example, Njirjak et al. report $40\%$ similarity among generated sequences, which is consistent with convergence toward a small set of high-fitness motifs. This concentration can limit exploration of more distant regions of sequence space and reduce the chance of uncovering alternative self-assembling motifs. Other generative models, as the previously mentioned VAEs and variants, can instead learn broader conditional distributions over sequences — yielding lower overall similarity while retaining controllability —, and are therefore promising approaches well suited to discover such alternative motifs.
@@
Crucially, aggregation propensity is only a first filter. Even when aggregation-prone sequences are identified, morphology remains a critical determinant of material function. Applications ranging from drug delivery to biosensors demand specific aggregate geometries (fibers, nets, spheres, etc.), yet datasets linking peptide sequence to computationally or experimentally validated self-assembly morphology are scarce. Recent efforts have started to steer material outcomes such as hydrogelation via human-in-the-loop approaches that couple experiments, simulation, and ML xuAcceleratingPredictionDiscovery2023, drazicUsingMachineLearning2025b; however, generative approaches that can reliably target specific supramolecular morphologies are yet to be explored, especially given that morphology annotations are fragmented across protocols and conditions. For example, efforts for data gathering concerning morphology have primarily assembled experimental databases such as SAPdb mathurSAPdbDatabaseShort2021, with simulation setups differing substantially across the aggregated studies. This lack of scale and consistency renders the available data unsuitable for training machine learning models. Without direct morphology labels, generative design must rely on indirect proxies as stand-ins to inject structural information; for example, predicted descriptors derived from the peptide sequence and the isolated monomer-level structure model. Features such as monomeric $\beta$-strand propensity, amphiphilicity, and charge distribution were postulated to directly influence whether peptides self-assemble into fibrils, sheets, or globular micelles wangAggregationRulesShort2024. For example, sequences dominated by hydrophobic residues may form compact, spherical aggregates (minimizing exposed hydrophobic surface). Tools like PEP-FOLD reyPEPFOLD4PHdependentForce2023 allow the effective generation of the most likely 3D conformations of isolated peptides in aqueous solution, which is one of the most common target solvents for supramolecular peptide assemblies.
@@
Computationally derived peptide-level descriptors may, therefore, serve as proxy signals that bias sequence discovery toward assembly behaviors of interest, even if they do not uniquely determine supramolecular morphology. To validate this hypothesis, we introduce PepMorph, a framework designed to condition peptide discovery on aggregation propensity and monomer–level descriptors for morphology awareness. At its core, PepMorph employs a transformer-based CVAE with a masking mechanism, akin to VAEAC, allowing peptide generation to be flexibly conditioned on any set of descriptors of isolated peptides that serve as morphology proxies. We demonstrate that the model generates highly novel and diverse peptides, with $10\%$ similarity, whose properties match all their targets in $55\%$ of cases. All the generated candidates are then passed through a dual-stage filtering pipeline, ensuring that only sequences meeting aggregation and sequence or structure requirements are retained. In end-to-end validation with CG-MD simulations, all sequences formed aggregates, and $83\%$ matched the intended morphology by visual inspection, under our CG-MD validation protocol and morphology criterion, supporting morphology-aware sequence discovery. PepMorph bridges the gap between unguided sequence exploration and fully constrained design, offering a versatile route toward developing functional supramolecular materials.
2 Results
@@
We begin by forming the PepMorph dataset, an extensive aggregation and isolated peptide descriptor dataset. Then, we implement and rigorously evaluate the PepMorph pipeline on penta-to-decapeptide morphology-aware aggregate discovery tasks. To do so, we train a transformer-based CVAE with a masking mechanism on an augmented dataset comprising aggregation metrics and computed descriptors for each peptide sequence. We then demonstrate the model’s generative performance by sampling sequences conditioned on assembly and structural metrics. Next, we apply our post-generation filters, selecting fifteen candidates each for fibrillar and spherical morphologies. Finally, we subject these peptides to CG-MD simulations in aqueous solution, quantifying their AP and examining the resulting aggregate morphologies quantitatively and visually.
@@
One of the main issues with predictive models for short peptides is the lack of extensive datasets like the Protein Data Bank, which provides a comprehensive and standardized structural reference for proteins zardeckiPDB101EducationalResources2022. Nevertheless, we can move towards machine learning models for assembly of short peptides by enriching existing datasets with additional sequences and morphology proxy features (Figure 1a).
@@
Figure 1: PepMorph dataset (a) Data curation and feature-extraction workflow: we merge three sources: Wang et al. wangDeepLearningEmpowers2023a ($\sim$62k peptides, 5-10 amino acids ($a$)), Teijlingen & Tuttle vanteijlingenTripeptidesTwoStepActive2021a ($\sim$60k, 3-8 $a$), and a set of $\sim$39k random peptides (retained set from successful PEP-FOLD runs, 5-10 $a$). After deduplication, self-assembly (SA) labels are assigned, and peptide conformations are predicted with PEP-FOLD for Wang et al. and random peptides to derive biophysical descriptors ($\beta$-strand assignment, net charge and hydrophobic moment). The resulting PepMorph corpus contains  161k unique peptides spanning 3-10 $a$ with aggregation-propensity (AP) values and SA/no-SA labels, as well the calculated peptide–level descriptors. Univariate summaries of the PepMorph dataset are shown, specifically of AP density (b), assembly vs no assembly (c), hydrophobic moment (d), peptide length (e), presence of $\beta$-strand assignment (f) and net-charge (g). Regions regarding no-assembly and assembly are highlighted in (b), and condition regions used when targeting specific morphologies are highlighted in the remaining summaries (d-g).
@@
We curated continuous AP values from Wang et al. wangDeepLearningEmpowers2023a and Teijlingen & Tuttle vanteijlingenTripeptidesTwoStepActive2021a, where AP follows the CG-MD-derived SASA ratio wangDeepLearningEmpowers2023a. Both studies used the MARTINI 2.2 force field. For the 56 peptides common to both datasets, we observed a mean AP difference of $0.089\pm 0.13$, supporting comparability, and thus mergeability, of the two AP sources (full statistical summary in Supplementary Material). During deduplication, we retained the values reported by Wang et al. to maintain internal consistency with the SA/no-SA labeling thresholds and downstream validation protocol adopted throughout this work. Subsequently, we assigned peptide a self-assembly/no-self-assembly (SA/no-SA) flag also based on AP following the operational labeling strategy of Wang et al. wangDeepLearningEmpowers2023a, liuEfficientPredictionPeptide2023a: the dataset is split into low-AP vs high-AP classes around the median AP, while excluding samples in a narrow band around the median to reduce ambiguity and label noise. Because the original study reports this procedure but not explicit numeric cutoffs, we reconstructed the corresponding thresholds from their released AP distribution, yielding $\mathrm{AP}\geq 1.8$ (SA) and $\mathrm{AP}\leq 1.65$ (no-SA), with intermediate values left unlabeled (Figure 1b). These thresholds should, therefore, be interpreted as operational separators rather than actual value–limiters for assembly. Applying this processing step yielded a merged dataset with AP for $121,652$ peptides, $93,668$ with a SA/no-SA label.
Given the masking mechanism of our model, further explained in the next section, PepMorph does not require any specific target to be present to generate sequences (e.g., AP can be absent). As such, we also added random sequences not present in the literature–derived set to expand the conditioning space. We sampled $50,000$ additional random sequences, by drawing the length uniformly from 3103-10 and sampling each residue independently from the 20 standard aminoacids — and avoiding duplicates from the Wang et al.’s AP data source. This also reduces compositional bias, as the original self-assembly–focused datasets are mildly enriched in aromatic residues, and adding uniformly sampled random sequences shifts the overall amino-acid frequencies closer to uniformity (Supplementary Figure 3). For this random set and for the subset of Wang et al.’s peptides that have an SA/no-SA label ($n=41,680$), we predicted isolated 3D conformations with PEP-FOLD 4 reyPEPFOLD4PHdependentForce2023, which would then permit the retrieval of the biophysical descriptors that can serve as proxies for aggregate morphology — computed in the isolated monomeric state and, thus, only used as conditioning signals, not direct aggregation labels. We restricted PEP-FOLD processing of the Wang et al.’s set to labeled peptides to focus descriptor extraction on sequences for which downstream assembly-related analyses are defined, and to manage the computational cost of structure prediction. In practice, PEP-FOLD is only applicable to peptides of length $\geq 5$; therefore, synthetic sequences of length 343-4 were excluded from 3D-descriptor extraction by construction. For the remaining lengths, PEP-FOLD almost always yields a usable lowest-energy model. Rare residual failures did still exist, but are negligible. Overall, PEP-FOLD succeeded for $41,402$ ($99.3\%$) SA labeled Wang peptides and for 39,468/50,000 ($78.9\%$) synthetic peptides, yielding 80,870 sequences with valid PEP-FOLD-derived descriptors.
As we will elaborate in the validation pipeline, we focus on three biophysical metrics that plausibly steer spherical versus fibrillar assembly:
@@
  • $\beta$-strand assignment, referring to whether or not any of the amino acid residues adopt an extended strand; peptides with $\beta$-strand propensity tend to stack via backbone hydrogen bonds into elongated, fibrillar (cylindrical) assemblies dehsorkhiSelfassemblingAmphiphilicPeptides2014.
@@
  • net charge, measuring the peptide’s overall charge at physiological pH; for example, highly charged peptides experience strong electrostatic repulsion that impedes self-assembly dehsorkhiSelfassemblingAmphiphilicPeptides2014.
We obtained valid 3D conformations (and thus retrievable descriptors) for $80,870$ sequences ($\approx 50\%$ of the full corpus). Importantly, this does not mean that the remaining sequences were discarded: PepMorph is trained on the full dataset of 161,120161{,}120 sequences, described in Table 1, with missing descriptor fields (including unavailable 3D descriptors) explicitly represented via the masking mechanism that is explained in the PepMorph Model Section. The distribution across sequence lengths is relatively balanced, with between $15,000$ and $35,000$ peptides represented at each length (Figure 1e). All possible tripeptides are included, but for longer lengths the dataset covers a smaller fraction of the exponentially expanding theoretical space of $20^{L}$ sequences (see Supplementary Material).
Table 1: PepMorph dataset composition (161,120 entries). It comprises the peptide sequence and length, two aggregation descriptors, and three morphology–proxy descriptors. Coverage counts non-null entries. Synthetic random peptides do not have AP/SA labels by construction; they contribute only sequence-derived fields and the 3D morphology–proxy descriptors. All records are retained for training; unavailable fields appear as missing values and are handled via masking.
@@
sequence Peptide in FASTA format (uppercase single-letter codes) zhanglabFASTAFormat wangDeepLearningEmpowers2023a, vanteijlingenTripeptidesTwoStepActive2021a + random $161,120$ ($100.0\%$)
length Number of residues in sequence derived
ap Aggregation propensity (AP) wangDeepLearningEmpowers2023a, vanteijlingenTripeptidesTwoStepActive2021a $121,652$ ($75.5\%$)
is_assembled Self-assembly label (1=SA, 0=no-SA) $93,668$ ($58.1\%$)
has_beta_strand Whether any residue is assigned to $\beta$-strand in the predicted peptide 3D conformation computed $80,870$ ($50.2\%$)
hydrophobic_moment Magnitude of hydrophobic moment from predicted peptide 3D conformation
net_charge Peptide net charge at neutral pH
The resulting PepMorph dataset also exhibits notable imbalances across descriptors. In particular, SA labels are roughly double that of no-SA (Figure 1c), and $\beta$-strand assignment is exceedingly rare (Figure 1f). The AP distribution itself is continuous, with many sequences clustering around the cutoff of $\geq\!1.8$ (Figure 1b). Adding to this, AP shows distinct distributions across lengths (Supplementary Figure 4), showing that the true cutoff can be length–dependent; even so, for simplicity, we keep this threshold across lengths. For these reasons, this threshold should be regarded as an approximate separator. By contrast, continuous descriptors such as the hydrophobic moment (Figure 1d) and net charge (Figure 1g) show a broad spread. However, most peptides lie in the low-to-moderate regime for the former. These patterns emphasize the potential bias and structural diversity in the PepMorph dataset. Its associated sparsity is, however, naturally handled by the aforementioned masking mechanism. Furthermore, extending the dataset with additional descriptors from the generated peptide conformations would also be trivial to integrate into the pipeline.
@@
To address this limitation and fully leverage our dataset, we develop a transformer-based CVAE that supports partial conditioning via an explicit masking mechanism, similarly to the arbitrary–conditioning paradigm of VAEAC — although in VAEAC this is used for generic feature imputation and image inpainting, rather than targeted masking of descriptors ivanovVariationalAutoencoderArbitrary2018 (Figure 2a).
Figure 2: PepMorph model and generation validation. (a) Schematic of the Transformer-based Conditional Variational Autoencoder with the masking mechanism: a descriptor vector $c$ and mask $m$ are summarized into a condition summary that conditions both the latent prior and the autoregressive Transformer decoder, enabling generation under arbitrary subsets of constraints. (b) Amino acid frequency in generated peptides closely follows the training distribution for both common and rare condition sets. (c) Novelty relative to the training set, quantified by the nearest-neighbour normalized edit distance (NED), showing the empirical cumulative fraction of conditioning targets whose mean NED to the closest training peptide is $\leq x$, which is similar in common vs. rare condition sets. (d) Condition-matching as a function of the number of conditioned descriptors $k$: the fraction of peptides meeting their targets declines as constraints tighten. (e) Similarity via Needleman-Wunsch percent identity of generated sequences (points) to the training set ($\mathrm{Sim}_{\text{train}}$) vs. generated sequences within the same common conditions set ($\mathrm{Sim}_{\text{gen}}^{\text{within}}$), color coded by the number of conditions $k$; values remain near low-identity baselines.